OpenStudy (anonymous):

Solve. X^2 - 6x + 4 = 0

OpenStudy (.cheesecake.):

Hi Welcome To Openstudy! :) Are we supposed to solve by factoring?

OpenStudy (just_one_last_goodbye):

@Kathcarl need help? :D

OpenStudy (anonymous):

Yes solve by factoring

OpenStudy (dayakar):

X^2-2*X*3 +3^2 - 3^2 +4 =0 [ making the equation into (a-b)^2 = a^2-2ab +b^2]

OpenStudy (dayakar):

(x-3)^2 -9 +4 =0 (x-3) ^2 -5=0 (x-3)^2 =5 can u solve remaining part

OpenStudy (anonymous):

Yes